### 30 Mar 2026 v1

# Debt Behind the AI Boom: A Large-Scale Empirical Study of AI-Generated Code in the Wild

Abstract.
AI coding assistants are now widely used in software development. Software developers increasingly integrate AI-generated code into their codebases to improve productivity. Prior studies have shown that AI-generated code may contain code quality issues under controlled settings. However, we still know little about the real-world impact of AI-generated code on software quality and maintenance after it is introduced into production repositories. In other words, it remains unclear whether such issues are quickly fixed or persist and accumulate over time as technical debt. In this paper, we conduct a large-scale empirical study on the technical debt introduced by AI coding assistants in the wild. To achieve that, we built a dataset of 304,362 verified AI-authored commits from 6,275 GitHub repositories, covering five widely used AI coding assistants. For each commit, we run static analysis before and after the change to precisely attribute which code smells, bugs, and security issues the AI introduced. We then track each introduced issue from the introducing commit to the latest repository revision to study its lifecycle. Our results show that we identified 484,606 distinct issues, and that code smells are by far the most common type, accounting for 89.1% of all issues. We also find that more than 15% of commits from every AI coding assistant introduce at least one issue, although the rates vary across tools. More importantly, 24.2% of tracked AI-introduced issues still survive at the latest revision of the repository. These findings show that AI-generated code can introduce long-term maintenance costs into real software projects and highlight the need for stronger quality assurance in AI-assisted development.
AI coding assistants, technical debt, code quality, software maintenance, empirical study
copyright: none
1. Introduction
Through AI coding assistants (e.g., Cursor, Claude Code), software developers can now describe what they want in natural language and get working code back in seconds. They significantly improve development productivity. Thus, AI is becoming standard equipment for modern software developers (GitHub, 2024). According to the 2025 Stack Overflow Developer Survey, 84% of professional developers are using or are planning to use AI tools within their development processes (Stack Overflow, 2025). The AI-generated code is also widely used in real-world software projects. For example, both Google and Microsoft disclosed in 2025 that AI now writes over 20% of their new code (Novet, 2025; Robison, 2024). Similarly, GitHub reported that more than 1.1 million public repositories used AI coding tools between 2024 and 2025  (GitHub, 2024). Overall, AI-generated code has graduated from experiment to production reality, and it is everywhere.
Although AI coding assistants have proven effective at generating functional programs, many previous research studies have revealed a range of quality concerns in AI-generated code. Recent studies have shown that AI-generated code suffers from functional bugs, runtime errors, and systemic maintainability issues (Liu et al., 2024; Siddiq et al., 2024). Also, the code produced by AI tools poses security risks (Pearce et al., 2025; Perry et al., 2023; Liu et al., 2025). Pearce et al. (Pearce et al., 2025) found that about 40% of AI-generated code in security-sensitive contexts contains critical vulnerabilities. However, recent research (Perry et al., 2023; Sabouri et al., 2025) has found that developers tend to place excessive trust in the quality of AI-generated code, blindly accepting the code without proper validation. As a result, these unverified code snippets are merged into production codebases. Over time, this can cause a considerable accumulation of technical debt, which can be costly and time-consuming to address (Harding and Kloster, 2024; Moreschini et al., 2026).
Recognizing these risks, recent empirical studies have started to investigate AI-generated code in real-world repositories. These studies cover a range of practical concerns, such as security weaknesses (Fu et al., 2025; Wang et al., 2025), project-level development velocity (He et al., 2025), pull request acceptance rates (Watanabe et al., 2025), and code redundancy (Huang et al., 2026). For example, He et al. (He et al., 2025) found that Cursor adoption in 807 GitHub repositories led to a transient velocity boost but persistent increases in code complexity. However, existing studies still have several limitations. First, most studies focus on a single tool or a narrow set of tools, which limits the generalizability of their findings. Second, many studies rely on proxy-based attribution (e.g., configuration files, self-disclosure signals, or classifiers), rather than directly tracking which code lines or files are AI-generated. Third, they mostly evaluate code at a single point in time, failing to capture the long-term lifecycle of the code. Thus, it remains unknown how AI-generated code actually ages in production. We still do not know whether the technical debt it introduces persists, gets refactored, or silently accumulates over time.
To bridge this gap, we conduct a large-scale empirical study to investigate the lifecycle of AI-introduced technical debt in the wild. Our approach consists of three steps. First, we build a large dataset of verified AI-authored commits. Instead of relying on classifiers or proxy signals, we use explicit Git metadata to identify commits generated by AI coding tools across thousands of GitHub repositories. Second, we perform a commit-level quality analysis. For each AI-authored commit, we run static analysis tools on the source code immediately before and after the change. This allows us to precisely identify which code smells, bugs, and security issues the AI introduced or fixed. Third, we conduct a debt lifecycle analysis. We track each introduced issue to the latest repository revision (HEAD) to determine whether it still survives or has been resolved. Through this design, our study provides the first comprehensive view of how AI-generated code ages in real-world software.
Contribution. To the best of our knowledge, this paper is the first to:
  • Conduct a large-scale empirical study of AI-introduced technical debt across five major AI coding assistants (i.e., GitHub Copilot, Claude, Cursor, Gemini, and Devin) and over 6,000 real-world GitHub repositories.
  • Perform commit-level differential analysis to precisely attribute code smells, bugs, and security issues to individual AI-authored commits.
  • Track the lifecycle of AI-introduced technical debt from the moment of introduction to the latest repository revision, revealing how it persists, evolves, or gets resolved over time.
Open Science. To support the open science initiative, we publish the studies dataset and a replication package, which is publicly available in GitHub. [^1] https://github.com/xxxx/xxxx
Paper Organization. Section 2 presents the background and motivation. Section 3 describes our approach. Section 4 presents the experimental setup. Section 5 presents the results. Section 6 discusses the findings and their implications. Section 7 presents the related work. Section 8 discloses the threats to validity. Section 9 draws the conclusions.
2. Background
2.1. AI Coding Assistants
Modern AI coding assistants (e.g., Cursor, Claude Code, GitHub Copilot) are now deeply embedded in software development workflows. They help software developers write, modify, explain, test, or debug code using natural-language instructions and code context. Advanced agentic tools can now process entire functions, files, or even entire repositories to autonomously create pull requests with minimal human intervention. Driven by these improved capabilities, AI-generated code is entering production codebases at unprecedented speed and scale. According to GitHub, over 1.1 million public repositories adopted AI coding tools between 2024 and 2025 (GitHub, 2024). As shown in Figure 1, in Anthropic’s claudes-c-compiler repository (Anthropic, 2026), Claude appears as the top contributor, with 3,957 commits and nearly 500K lines of code added within just a few weeks. At this volume and speed, it is unlikely that all AI-generated code receives a thorough human review. This makes it increasingly important to understand the long-term implications of AI-generated code on software quality and maintenance.
Figure 1. Contributor statistics for the Anthropic claudes-c-compiler repository.
In controlled (lab) settings (Pearce et al., 2025; Liu et al., 2024), the provenance of code is usually clear since researchers can generate the code directly. However, it is different in real-world codebases. The reason is that AI-generated code and human-written code are often mixed together. This makes it harder to observe the long-term impact of AI-generated code after it is merged into production codebases.
$2.$ Technical Debt
Technical debt refers to design or implementation choices that prioritize short-term speed over long-term quality (Cunningham, 1992). These shortcuts may help in the short term, but they increase the future cost of maintaining and evolving the software (Avgeriou et al., 2016; Li et al., 2015). Its costs can accumulate over time if not addressed. This concern becomes more important as AI coding assistants are widely adopted. AI coding assistants help developers work faster and produce more code. Previous studies have shown that AI-generated code contains code smells, maintainability issues, and security vulnerabilities (Liu et al., 2024; Pearce et al., 2025; Fu et al., 2025). When such issues are accepted into production repositories, they can accumulate as technical debt in the codebase.
seagullz4/hysteria2 (seagullz4, 2025c), hysteria2.py:L77, e277daf $\Downarrow$ seagullz4/hysteria2, hysteria2.py:L77 d9e392d
Figure 2. Command injection risk introduced by GitHub Copilot in hysteria2 (1.7K stars) and its subsequent fix.
realsenseai/librealsense (RealSense, 2025), test-fps-performance.py:L168, 5535b8a $\Downarrow$ realsenseai/librealsense, test-fps-performance.py:L37, 14026c8
Figure 3. Undefined variables introduced by GitHub Copilot in librealsense (8.6K stars), causing a runtime error. Fixed three weeks later.
2.3. Motivation
The technical debt introduced by AI coding assistants is not just a theoretical risk. Such issues are becoming increasingly common in real-world codebases. Figure 2 and 3 show two examples. In Figure 2, a Copilot-authored commit introduced a shell=True subprocess call in hysteria2.py (seagullz4, 2025b). This pattern increases security risk by allowing command injection if user input is involved. A human developer later fixed it, noting in the commit message: “Improve code security by removing shell=True from subprocess calls” (seagullz4, 2025a). Figure 3 shows a bug from Intel’s librealsense (RealSense, 2025). A Copilot-authored commit replaced a value with a named constant but never defined it (Intel RealSense, 2025b). The buggy code remained in the repository for over three weeks before the maintainer added the missing definition (Intel RealSense, 2025a).
These two examples highlight the motivation of our study. AI coding assistants can generate functional code, but they may introduce quality issues into production codebases. Developers also tend to over-trust and accept AI suggestions without thorough review (Perry et al., 2023; Sabouri et al., 2025). These issues may be fixed later, or they may persist for a long time, or even create long-term maintenance challenges. Recent studies (Nguyen and Nadi, 2022; Liu et al., 2024; Fu et al., 2025; Siddiq et al., 2024; Wang et al., 2025; He et al., 2025) have examined the AI-generated code, but they have several limitations. However, most prior studies focus on a single tool, a small set of tasks, or controlled settings. For example, Watanabe et al. (Watanabe et al., 2025) measured the initial acceptance rate of AI-authored code in a single repository. We still know little about the long-term implications of AI-generated code on software quality and maintenance. Thus, we study the technical debt introduced by AI coding assistants in the wild: how often it is introduced, what kinds of issues appear, and whether those issues still remain in the codebase over time.
3. Approach
Figure 4. Overview of our approach.
Figure 4 provides an overview of our approach. We first collect AI-authored commits from GitHub repositories at scale (Section 3.1). We then analyze each AI-authored commit at the code level to determine which quality issues it introduced or fixed (Section 3.2). Finally, we track the lifecycle of both the issues and the code itself to determine whether AI-introduced debt persists or gets resolved over time (Section 3.3).
3.1. Data Collection
This step aims to identify candidate GitHub repositories that contain AI-authored commits.
Repository Discovery.  We use the GitHub Archive dataset, which records all public events (i.e., PushEvent) on GitHub, to identify repositories with potential AI-authored code. We scan all PushEvent records from January 2024 to October 2025, focusing on repositories with recent development activity. Using Google BigQuery, we extract four metadata fields from each event: actor login, author name, author email, and commit message. We then match these fields against our curated AI-attribution patterns. Only repositories with at least one matching event are retained. It is worth mentioning that GitHub changed the Events API on October 7, 2025, and removed commit-level summaries from PushEvent payloads (GitHub, 2025). Thus, for the post-October-2025 data, we supplement BigQuery-based discovery with GitHub API queries over top repositories and then apply the same attribution rules through full-history repository scanning (described in the next section). These steps help us ensure that active and popular repositories are not missed due to the API change.
AI Attribution Rules.  We build attribution rules for widely adopted AI coding tools (e.g., Cursor, GitHub Copilot, Claude Code) identified in the 2025 Stack Overflow Developer Survey (Stack Overflow, 2025). We identify AI-authored commits using explicit signals in Git metadata. Our approach covers AI-authored commits only when the use of an AI coding tool leaves explicit traces in Git metadata The rules are based on four sources of evidence: (1)  actor logins (e.g., copilot-swe-agent[bot]), (2) author emails (e.g., noreply@anthropic.com), (3) author names (e.g., Cursor Agent), and (4) Co-authored-by trailers in commit messages. These signals identify whether and which AI coding tool was involved in a commit. In total, 29 AI coding tools left identifiable traces in the repositories we collected. The full list of tools and detection rules is included in our replication package.
Full-History Commit Scanning.  The discovery stage captures only push-event metadata. However, it provides only partial evidence about AI-authored activity in a repository. To obtain a more complete set of AI-authored commits, we perform a bare clone of each candidate repository. We then scan the full commit history across all branches and apply the same attribution rules to every commit. For each commit, we extract the SHA, author and committer metadata, timestamp, and full commit message. This step allows us to identify AI-authored commits that are not directly visible during repository discovery (e.g., commits on non-default branches, commits outside the observation window).
Filtering.  To focus on established open-source projects, we filter out repositories that do not meet our study criteria. We keep only repositories with at least 100 GitHub stars. We remove forks with fewer than 100 stars. We also require at least one confirmed AI-authored commit. Our downstream analysis is restricted to production Python, JavaScript, and TypeScript source files, excluding repositories or commits that do not contain analyzable source code in these languages. After filtering, our dataset contains 6,275 repositories with 304.4k total AI-authored commits.
3.2. Commit-Level Quality Analysis
To understand the impact of AI coding tools, we avoid evaluating a repository snapshot at a single point in time. For each AI-authored commit $c$, we analyze the source code at its parent revision (before) and at the commit itself (after) to determine which quality issues it introduced or fixed. Our analysis focuses on production source files written in Python, JavaScript, and TypeScript. We exclude files that are unlikely to reflect production code quality (e.g., tests, documentation, configuration files, auto-built artifacts, and vendored dependencies). Files are classified based on their paths and naming patterns (e.g., files under test/ or __tests__/ directories, or matching *_test.py). The full classification rules are included in our replication package.
Static Analysis.  For each AI-authored commit $c$ that modifies a source file $f$, we check out two versions of $f$ (i.e., the version before $c$ (at its parent commit) and the version after $c$). We run the same static analysis toolchain on both versions to identify code quality and security issues. For Python, we use Pylint and Bandit. For JavaScript and TypeScript, we use ESLint and njsscan. For each detected issue, we record its rule identifier, line number, severity, and message. This produces two issue sets II: the issue set before the commit, denoted by $I_{f}^{-}$, and the issue set after the commit, denoted by $I_{f}^{+}$.
Differential Attribution.  To find out which issues commit $c$ introduced or fixed, we compare $I_{f}^{-}$ and $I_{f}^{+}$. We also use git diff to extract the set of changed lines, denoted by $\Delta_{f}$. However, not all differences between $I_{f}^{-}$ and $I_{f}^{+}$ (i.e., issues in $I_{f}^{+}\setminus I_{f}^{-}$ or $I_{f}^{-}\setminus I_{f}^{+}$) represent real changes. When a commit inserts or deletes lines, it can cause existing issues to shift line numbers. To address this, we first match issues across the two sets. An issue $i$ is considered matched if the same rule and message appear in both $I_{f}^{-}$ and $I_{f}^{+}$ at the same or nearby line number. After matching, the remaining issues are classified as follows. An unmatched issue $i\in I_{f}^{+}$ is classified as introduced only if its line falls within $\Delta_{f}$. This means the issue exists only after the commit, on a line that the commit actually changed. An unmatched issue $i\in I_{f}^{-}$ is classified as fixed. This means the issue existed before the commit but is no longer present afterward.
$3.$ Debt Lifecycle Analysis
Detecting technical debt at the time of introduction is only half the picture. An issue that is quickly resolved has a very different cost than one that lingers for months. We therefore track whether AI-introduced issues persist or get resolved over time.
Issue Survival.  For each issue introduced by an AI-authored commit, we check whether it still exists at the repository’s latest revision (i.e., HEAD). If the file has been renamed, we follow its history using git log --follow. We then run static analysis on the corresponding file at HEAD. Next, we look for the same issue in the analysis results. We do not rely on the line number alone, since the location of the issue may move as the file changes. Instead, we match issues using their rule identifier together with a small amount of surrounding code context. If a match is found, the issue is classified as surviving. Otherwise, it is classified as not surviving.
At the same time, we also record whether files touched by AI-authored commits are modified again before HEAD. We trace the subsequent commit history of each affected file to understand how actively it is maintained after the AI-authored change. This additional context helps us interpret the survival results and understand the maintenance patterns around AI-introduced debt.
4. Experimental Setup
In this section, we describe the experimental setup for our study.
Table 1. Summary of AI-authored commits by coding tool.
AI Coding Tool # AI Commits # Repos Avg. Commits/Repo
GitHub Copilot 117,851 4,473 26.3
Claude 139,300 2,431 57.3
Cursor 19791 717 27.6
Gemini 12770 681 18.8
Devin 14650 300 48.8
Total 304,362 6,275* 48.5
*Repositories may use more than one tool; the total is deduplicated.
4.1. Dataset Summary
We collected our dataset by mining GitHub and applying filtering criteria. In our study, we focus on repositories containing production source files in Python, JavaScript, or TypeScript. Also, each repository must have at least one AI-authored commit. After filtering, we found 6,699 GitHub repositories with AI-authored commits, covering 29 AI coding tools. However, some tools have few commits, which may not provide reliable data for comparison. Thus, we focus on the five assistants with more than 10,000 attributed commits: GitHub Copilot, Claude, Cursor, Gemini, and Devin. Our final dataset includes 6,275 public GitHub repositories with 304,362 verified AI-authored commits. Table 1 summarizes the distribution of commits across the five AI coding assistants in our dataset. Figure 5(a) shows the monthly growth of AI-authored commits, with a sharp increase starting from mid-2025. Figure 5(b) shows that our dataset covers repositories with a wide range of popularity levels.
(a) Monthly commits.
(b) Repository distribution.
Figure 5. Overview of our dataset: (a) growth of AI-authored commits over time, and (b) distribution of repositories by GitHub star count as of March 2026.
4.2. Research Questions
Our study focuses on the following research questions:
RQ1: What kinds of technical debt are introduced by AI coding assistants? This question investigates the basic characteristics of AI-introduced technical debt. We study what types of debt they introduce, including code smells and security issues. We also analyze how these issues are distributed across severity levels, languages, and rules.
RQ2: How does technical debt vary across AI coding assistants? This question compares different AI coding assistants at the commit level. We examine whether some assistants introduce more technical debt than others, and whether the kinds of issues they introduce differ. This helps us understand whether technical debt patterns are tool-specific.
RQ3: How does AI-introduced technical debt evolve over time? Introducing debt is not necessarily a problem if it gets fixed quickly. The real concern is debt that persists unnoticed. We study whether introduced issues remain in the latest version of the repository or disappear over time. We also examine whether the associated code and files are later modified, fixed, or refactored.
4.3. Evaluation Metrics
We use the following metrics to support our three research questions. For issue introduction (RQ1, RQ2), we report the total number of issues introduced by AI-authored commits, the percentage of commits that introduce at least one issue, and the average number of issues per commit. We break these down by issue type, programming language, rule, and AI coding assistant.
For the debt lifecycle (RQ3), we use three metrics. First, we compute the net impact by comparing the number of issues introduced and fixed by AI-authored commits. Second, we measure the survival rate of introduced issues:
$\textit{Survival Rate}=\frac{\text{\# issues surviving at {HEAD}}}{\text{\# issues tracked}}$
Third, to normalize for differences in commit volume across time periods, we report the number of surviving issues per 100 AI-authored commits.
5. Results
This section presents the empirical results of our study and answers the three research questions raised in Section 4.2.
Table 2. Overview of AI-introduced technical debt by issue type.
Type # Introduced # Repos # Commits % of Total
Code Smells 431,850 3,789 25,895 89.1%
Runtime Bugs 28,149 663 1,646 5.8%
Security Issues 24,607 1,038 4,158 5.1%
All Issues 484,606 3,841 26,564 -
5.1. RQ1: Types and Patterns of AI-Introduced Debt
Overview.  Table 2 presents a summary of the technical debt introduced by AI coding assistants in our dataset. In total, we identified 484,606 introduced issues across 3,841 repositories (61.2% of 6,275 repositories) and 26,564 commits (8.7% of 304,362 commits). This shows that a non-trivial portion of AI-authored commits introduce quality issues, and that these issues affect a large number of real-world repositories. Among all introduced issues, code smells, runtime bugs, and security issues are the three main categories. Code smells are by far the most common, accounting for 89.1% of all introduced issues. Below, we discuss each type in detail with real-world examples.
Table 3. Top 5 most frequent rules violated by AI coding assistants for each issue type.
Type Rule Count Rate
Code Smells Broad exception handling 41,723 8.6%
Unused variables or parameters 28,718 5.9%
Unused argument 24,444 5.0%
Shadowed outer variable 20,251 4.2%
Access to protected member 19,835 4.1%
Runtime Bugs Undefined variable or reference 23,091 4.8%
Redeclared symbol 1,870 0.4%
Possibly used before assignment 1584 0.3%
Access member before definition 893 0.2%
Unsubscriptable object 130 0.0%
Security Issues Subprocess Without Shell Check 4,334 0.9%
Try-Except-Pass 4,040 0.8%
Partial Executable Path 2,539 0.5%
Insecure Random Generator 2,355 0.5%
Subprocess Import 1,718 0.4%
ArchiveBox/ArchiveBox (archivebox, 2025c), core/models.py:L594:598, d360798
Listing 1: Code smell: broad exception handling and missing file encoding in ArchiveBox (Claude Code).
Code Smells.  Code smells are maintainability problems that make code harder to understand, debug, and evolve (Fowler, 2018). They increase long-term maintenance costs, even if they do not cause immediate failures. This finding is consistent with prior work under controlled settings (Liu et al., 2024; Siddiq et al., 2024), but our study confirms that the same pattern also appears in real-world repositories. Table 3 lists the top 5 most common code smell patterns (e.g., broad exception handling, unused variables or parameters). These issues are often small and easy to overlook during code review. Listing 1 shows an example from ArchiveBox (archivebox, 2025a) (>>27k stars). In commit d36079829bed (archivebox, 2025c), Claude Code updated the metadata loading logic in ArchiveBox. But the new code introduces two code smells. First, the bare except: pass block catches all exceptions silently. This makes errors harder to detect and debug. Second, the open() function does not specify a file encoding. This can lead to inconsistent behavior across different platforms and locales, since the default encoding may vary (Naoki, 2021). These issues may not cause immediate failures, but they can lead to maintenance challenges and subtle bugs in the future.
firecrawl/firecrawl (firecrawl, 2025b), python-sdk/firecrawl/firecrawl.py:L4004:4047, fb99747
Listing 2: Runtime bug: undefined variable causing NameError in Firecrawl (Devin).
Runtime Bugs.  Runtime bugs are code defects that can cause the program to fail during execution. Compared with code smells, they are less frequent. From Table 2, 28,149 runtime bugs are identified, which cover 663 repositories and 1,646 commits. However, their impact is more direct and severe than code smells. Table 3 shows the top 5 most common runtime bugs, which include undefined variable or reference, redeclared symbol, access to member before definition, possibly used before assignment, and unsubscriptable object. These patterns suggest that AI-generated code may look locally correct, but still fail to stay consistent with the surrounding context. What is interesting in this table is that we identified 23,091 cases of undefined variable or reference. Listing 2 presents one such case from firecrawl (>>98k stars). In commit fb99747ba978 (firecrawl, 2025b), Devin added a call that passes cache=cache as an argument. However, cache is never defined in the method, which leads to a NameError when that path is executed. The maintainer later fixed the bug by removing the undefined argument (firecrawl, 2025a). This example shows that AI-generated code can introduce real runtime errors. These errors require additional human effort to fix later.
microsoft/data-formulator (Microsoft, 2025), /tables_routes.py:L881:889, d8549c0
Listing 3: Security issue: possible SQL injection in data-formulator (Copilot).
Security Issues.  Security issues are another concern in AI-generated code. They pose a threat to data privacy and system safety. If not properly addressed, they can lead to data breaches or unauthorized access. Thus, it is important to identify and fix security issues early before they enter production repositories. As shown in Table 2, potentially insecure code patterns are also detected in 1,038 repositories and 4,158 commits. Table 3 shows that the most common security issues include subprocess without shell check, try-except-pass, insecure random generator, partial executable path, and unsafe HuggingFace Hub download. These patterns suggest that AI-generated code can introduce unsafe practices in process execution, error handling, and external resource usage. Figure 2 shows an example of the most common issues (i.e., subprocess without shell check) from hysteria2 (>>1.5k stars). A Copilot-authored commit (seagullz4, 2025b) introduced a shell=True subprocess call, which can enable command injection if user input reaches the call. A human developer later identified and removed the unsafe flag (seagullz4, 2025a). Beyond that, Listing 3 shows another example of a security issue (SQL injection) in data-formulator (>>1.2k stars). This repository is developed by Microsoft and uses GitHub Copilot for code generation. In commit d8549c0 (Microsoft, 2025), Copilot added a backend endpoint that constructs a SQL query by directly interpolating a user-supplied table name. This creates a potential SQL injection vector. If an attacker can control the source_name variable, they can inject malicious SQL code that may lead to data breaches or unauthorized access. This issue remained in the repository for several weeks before the maintainer refactored the code and removed the unsafe SQL construction (Microsoft, 2026).
Table 4. Top 5 most frequent rules by language. Rate indicates each rule’s share of all issues in that language.
Language Rule Count Rate
Python Broad exception handling 41,721 14.1%
Unused argument 24,444 8.3%
Undefined variable or reference 23,091 7.8%
Access to protected member 19,835 6.7%
Unused import 17,435 5.9%
JavaScript/ TypeScript Unused variables or parameters 28,718 15.2%
Shadowed outer variable 18,016 9.6%
Block-scoped variable misuse 11255 6.0%
No sequences $89$ 4.8%
No unused expressions 7673 4.1%
Programming Language Differences.  Table 4 compares the top 5 most frequent rules in Python and JavaScript/TypeScript. There are some patterns that are common in both languages. For example, both languages show issues related to unused code (i.e., unused arguments in Python, unused variables in JavaScript/TypeScript). At the same time, each language also has its own characteristic issues. Python’s top rules are dominated by exception handling and dynamic typing problems, while JavaScript/TypeScript issues tend to involve scoping and variable declaration patterns. This observation is consistent with prior studies (Liu et al., 2024; Siddiq et al., 2024), which also found that the types of issues in AI-generated code can vary depending on the programming language and the tools used. These results suggest that some debt patterns may be language-specific. However, the overall trend (e.g., code smells dominate) holds across both languages.
Answer to RQ1: AI-generated code introduces technical debt in the form of code smells (89.1%), runtime bugs (5.8%), and security issues (5.1%). Among them, code smells are by far the most common.
Figure 6. Percentage of commits with issues and total commit volume per AI coding assistant.
Table 5. Average number of issues introduced per commit by type. Darker cells indicate higher values.
5.2. RQ2: Comparison Across AI Coding Assistants
In this RQ, we examine how technical debt patterns vary across AI coding assistants. Figure 6 shows the percentage of commits with issues for each of the five AI coding assistants. What stands out in this figure is that more than 15% of commits by each AI tool introduce at least one issue. The rates also vary across tools, ranging from 17.3% for GitHub Copilot to 28.7% for Gemini. However, technical debt does not disappear when using different or newer tools. This suggests that technical debt is a common and persistent problem in AI-generated code.
Table 5 further compares the average number of introduced issues per commit by type. We can see that all five tools share a common pattern, where the code smell rate is much higher than the bug and security issue rates. This is consistent with our findings in RQ1. At the same time, there are also differences across tools. From Table 5, we can see that Claude has the highest issue rate per commit (1.96), while Devin has the lowest (0.87). These differences may be due to differences in usage patterns and development context, rather than the tools alone. Still, it is apparent that the overall pattern of technical debt is consistent across all five tools.
Answer to RQ2: Technical debt patterns vary across AI coding assistants, but the overall trend is consistent. Across all five tools, code smells remain the dominant form of AI-introduced debt.
Figure 7. Net impact of AI coding assistants: issues introduced vs. fixed by issue type.
5.3. RQ3: Lifecycle of AI-Introduced Debt
Net Impact.  In RQ1 and RQ2, we focus on the technical debt introduced by AI-authored commits. However, AI coding assistants can also remove existing issues during refactoring or code improvement. To better understand the overall lifecycle of AI-introduced debt, we compare the number of issues introduced and fixed by AI commits (see Figure 7). For code smells, we can see that AI-authored commits fix more issues than they introduce (449,984 vs. 431,850), resulting in a net reduction of 18,134 code smells. In contrast, for runtime bugs and security issues, AI commits introduce more issues than they fix. What is interesting is that AI introduces nearly twice as many security issues as it fixes. These findings indicate that the net impact of AI coding assistants is mixed. AI coding assistants can help reduce maintainability issues, which tend to follow simple and repetitive patterns. However, for bugs and security issues, which require a deeper understanding and reasoning about program logic and context, AI coding assistants introduce more problems than they resolve.
Figure 8. Cumulative growth of AI-introduced issues over time, by issue type.
Table 6. Issue survival by time since introduction. Values show survival rates and the number of surviving issues per 100 AI-authored commits.
Time Since Introduction Survival Rate Surviving Issues per 100 Commits
Overall Code Smells Bugs Security
>> 9 months 19.20% 22.20 19.58 0.95 1.67
6–9 months 17.90% 29.73 25.60 1.56 2.57
3–6 months 29.30% 41.23 35.62 2.62 2.99
<< 3 months 24.80% 39.92 33.03 3.10 3.79
All Issues 24.20% 37.25 31.49 2.55 3.22
Issue Survival.  The net impact analysis above provides an overview of what AI coding assistants add and remove. But it does not show what happens to the specific issues introduced by AI. To answer this question, we track each AI-introduced issue to the latest repository snapshot and check whether it still exists at HEAD. Figure 8 shows that the cumulative number of surviving issues keeps growing over time. The total volume of unresolved technical debt increases rapidly, climbing from just a few hundred issues in early 2025 to over 110,000 surviving issues by February 2026. This suggests that as the rapid adoption of AI coding assistants continues, the amount of AI-introduced debt in real-world repositories is also growing significantly.
Table 6 provides a normalized view of this surviving debt. Overall, we can see that 24.2% of tracked AI-introduced issues potentially still survive at HEAD, which corresponds to 37.25 surviving issues per 100 AI-authored commits. Survival rates also differ across issue types. Security issues are the most likely to remain at HEAD (41.1%), followed by runtime bugs (30.3%) and code smells (22.7%). When we group these issues by age, we can see an interesting pattern. It is apparent that issues introduced earlier tend to have fewer survivors per 100 commits than more recent ones. For issues older than 9 months, only 22.20 issues per 100 commits still survive, compared to 39.92 for those introduced in the last 3 months. A similar decrease can also be observed for all three issue types. This suggests that AI-introduced debt does get gradually resolved over time. However, the resolution process is slow, and even after nine months, a substantial amount of debt remains in the codebase.
Figure 9. A TypeScript lint issue introduced by a Claude-authored commit in Stirling-PDF was fixed one day later by removing the unused variable.
These aggregate results are also reflected in real-world repositories. For example, the broad exception handling issue in Listing 1 was fixed within hours after it was introduced (archivebox, 2025c, b). Similarly, in Stirling-PDF (Tools, 2026) (>>75k stars), a Claude-authored commit introduced an unused variable filename in a TypeScript file (Tools, 2025b). As shown in Figure 9, the maintainer fixed it the next day with a commit titled “Fix TypeScript linting error” (Tools, 2025a). In contrast, the undefined variable bug in firecrawl (Listing 2) took 42 days before the maintainer fixed it (firecrawl, 2025a). However, some issues can survive for much longer or even remain unresolved. For instance, a Devin-authored commit in brave_search_tool.py added a call to requests.get(...) without a timeout on December 2024 (crewAIInc, 2024). This is a known potential security issue, because requests made without a timeout may block indefinitely if the remote service does not respond(PyCQA, 2023). However, the issue still remains in the latest repository revision.
Answer to RQ3: AI-introduced debt is not always removed quickly after introduction. Overall, 24.2% of tracked issues still survive at HEAD, and even issues older than 9 months leave a substantial surviving burden in the codebase.
6. Discussion
6.1. Implications
Our empirical study shows that AI coding assistants introduce technical debt into real software repositories. This is not a property of any single tool. We observed that across all five tools we studied, more than 15% of commits introduce at least one detectable issue. These issues persist regardless of repository size or popularity. In this section, we discuss the implications of our findings for practitioners, researchers, and tool builders.
The main implication is not just that AI coding assistants may produce low-quality code. More importantly, AI-assisted software development changes how technical debt enters and remains in production systems. Prior studies have shown that developers are more likely to over-trust AI suggestions (Sabouri et al., 2025). This over-trust can lead to a higher acceptance rate of AI-generated code, even when it contains issues. It means that many AI-introduced issues can accumulate in the codebase. Our results show that code smells are the most common type of AI-introduced debt. They often do not break the software system immediately, making them easy to accept during code review. But AI coding assistants allow developers to produce code at much higher speed and volume. As a result, these minor issues can accumulate into a substantial maintenance burden over time. Figure 8 shows that the cumulative number of surviving AI-introduced issues continues to rise over time, exceeding 110,000 by February 2026. The technical debt introduced by AI does not have to be a temporary side effect. It can become part of long-term maintenance challenges for modern software systems.
At the same time, our findings suggest although AI coding assistants introduce technical debt, they also fix existing issues in the codebase. First, we observe that AI co-authored commits actually fix a similar number of code smell issues as they introduce. This suggests that AI coding assistants are able to perform local cleanup and repetitive maintenance tasks effectively. They can recognize and address surface-level code quality problems (e.g., formatting, naming, or simple refactoring opportunities). However, what is concerning is that AI coding assistants seem to be less effective at fixing bugs and security issues, and they even introduce more of these than they fix. Section 5.3 shows that the survival rate of AI-introduced runtime bugs and security issues is much higher than that of code smells. This inconsistency suggests that AI coding assistants may struggle with changes that require deeper understanding of program behavior, execution context, or security implications. Also, the practical impact of bugs and security issues is often more severe than that of code smells, making it more critical to address them effectively. Developers should not treat all AI-generated code as equally trustworthy, and they should pay particular attention to changes that may introduce bugs or security vulnerabilities.
Our cross-tool comparison shows that this problem cannot be solved simply by switching from one assistant to another. Figure 6 shows that all five tools introduce a similar pattern of issues. They all have a high rate of code smells, and a non-trivial rate of bugs and security issues. This suggests that the quality risk is a systemic issue with the current mode of AI-assisted development. Thus, developers and teams should carefully review AI-generated code regardless of the tool used. Static analysis, tests, and security checks should be part of the normal workflow. Code review should extend beyond the point of merge. The main reason is that our study shows that 24.2% of tracked AI-introduced issues still survive at HEAD, and older issues are not fully cleaned up even after months. Merging the AI-generated code does not mean the end of the story since debt can persist and accumulate over time. This makes continuous monitoring and targeted debt repayment necessary for AI-touched code.
Overall, our results suggest that future research and tool design should not just focus on generating more acceptable code. We also need to ask whether that code remains maintainable, correct, and secure over time. Most existing research studies (Liu et al., 2024; Siddiq et al., 2024) focus on short-term outcomes such as task completion, acceptance rate, or immediate correctness. However, these measures only capture what happens when the code is introduced, not what happens later in maintenance. Future work needs to examine which factors make AI-introduced debt more likely to persist (e.g., repository maturity, review intensity, or task type). Our findings also suggest the need for better assistants. Future tools should make stronger checks for security-sensitive changes, use repository context more effectively, and clearly show AI provenance so reviewers can better judge the risk of a change. In the end, the key question is not only whether AI can produce code at scale, but whether the software engineering ecosystem can manage that code well over time.
6.2. Limitations
Lack of human-written baseline.  Our study does not compare AI-authored commits against a clean baseline of human-written commits. In real world, developers may use AI coding assistants but they do not leave any trace of it in the Git metadata. Even when a commit is labeled as AI-authored, the final code may include contributions from the human developer (e.g. edits, additions). This makes it difficult to construct a reliable baseline of purely human-authored commits. Further, comparing our datasets with an unreliable baseline would cause biased results and misleading conclusions. For this reason, we focus on tracking the technical debt inside explicitly confirmed AI-authored or co-authored commits.
Scope of technical debt.  In this study, we focus on code-level technical debt that can be detected by static analysis tools. We categorize issues into three types: code smells, bugs, and security issues. This categorization is based on the capabilities of the tools we use (e.g., Pylint, Bandit) and the common types of technical debt in software development. However, technical debt is a broader concept that includes architectural debt, design debt, documentation debt, and other forms of maintenance challenges. Thus, our findings should be interpreted as reflecting the code-level technical debt introduced by AI coding assistants, rather than the full spectrum of technical debt that may arise in software projects.
Survival Tracking Limitations.  Our issue survival analysis determines whether an issue still exists at HEAD by matching rule identifiers and surrounding code context. However, a file may be deleted or entirely rewritten between the introducing commit and HEAD. In such cases, the issue is classified as resolved, even though the resolution may not be a deliberate fix. We acknowledge this limitation and note that our survival rates may slightly underestimate the true persistence of AI-introduced debt.
7. Related work
Code Quality of AI-Generated Code.  Previous studies have examined the quality and security of AI-generated code. Based on controlled experiments, they showed that AI coding assistants (e.g., GitHub Copilot and ChatGPT) are able to produce functional code, but the code quality varies widely across languages, tasks, and prompts (Nguyen and Nadi, 2022; Liu et al., 2024; Mastropaolo et al., 2023b). AI-generated code can also contain security weaknesses, and developers may over-trust it and fail to properly review it (Pearce et al., 2025; Perry et al., 2023; Sandoval et al., 2023). Recent studies have also begun examining AI-generated code in real-world production environments. They show that AI-generated code is being widely adopted in platforms such as GitHub and Stack Overflow, and that it can carry quality and security issues (Fu et al., 2025; Siddiq et al., 2024; Wang et al., 2025). He et al. (He et al., 2025) studied the impact of Cursor adoption on 807 repositories and observed persistent increases in code complexity. Watanabe et al. (Watanabe et al., 2025) found that most Claude Code pull requests are merged, though many require human revisions. However, these studies mainly focus on a single tool or a narrow set of quality issues, and they do not track how those issues evolve.
Technical Debt in Software Development.  Technical debt is a core topic in software engineering research for years. Previous work has introduced many automated methods to identify code smells and self-admitted technical debt in software repositories (Yan et al., 2018; Ren et al., 2019; Palomba et al., 2018). Beyond detection, many studies have also investigated the lifecycle of technical debt in human-written code. Tufano et al. (Tufano et al., 2017) showed that code smells are often introduced during normal development. Once introduced, they can linger in the codebase for a long time before anyone removes them (Tufano et al., 2017). Digkas et al. (Digkas et al., 2018) identified a similar pattern in large open-source ecosystems. Their results showed that technical debt mainly accumulates when new code is added (Digkas et al., 2018, 2020). Other studies further suggest that debt is rarely removed in an intentional way, and that automated repayment is still difficult in practice (Zampetti et al., 2018; Mastropaolo et al., 2023a). Together, these studies provide a strong foundation for understanding technical debt in human-written software.
8. Threats to Validity
Below, we discuss threats that may impact the results of our study.
External validity. Our study focuses on public GitHub repositories with at least 100 stars and production source files in Python, JavaScript, and TypeScript. Therefore, our findings may not generalize to private repositories, smaller projects, or software written in other languages. In addition, we only analyze AI-authored commits that leave explicit traces in Git metadata. Projects or tools that use AI without such traces are outside the scope of our dataset.
Internal validity. Our pipeline depends on the correctness of AI attribution, issue matching, and lifecycle tracking. Although we use explicit Git metadata rather than proxy signals, some commits may still include both AI and human contributions. Similarly, matching issues across revisions is challenging because files evolve over time and issues may shift location. To reduce this risk, we compare code before and after each commit, restrict introduced issues to changed lines, and use rule identifiers, messages, and surrounding code context during matching. However, some attribution errors may still remain.
Construct validity. Technical debt is a broad concept with many possible forms. In this study, we operationalize technical debt mainly through code smells, correctness issues, and security issues detected by static analysis tools. This choice allows us to measure debt consistently at scale and track it over time at the commit level. However, static analysis tools can produce false positives, flagging code that is technically correct but matches a known risky pattern. Additionally, static analysis does not cover all forms of technical debt. Architectural debt, design erosion, documentation debt, and test adequacy issues are outside the scope of our tools. Our results should therefore be interpreted as evidence about code-level technical debt, not the full spectrum of quality challenges that AI-generated code may introduce.
9. Conclusion
AI coding assistants are rapidly becoming part of real-world software development, but their long-term impact on software quality remains unclear. In this paper, we presented a large-scale empirical study of the technical debt introduced by AI-generated code in the wild. By mining 304,362 AI-authored commits from 6,275 GitHub repositories, we designed a commit-level pipeline to identify introduced technical debt and track its later evolution in production repositories. Our study provides a longitudinal view of how AI-generated code ages after it is merged, including what kinds of debt are introduced, how debt varies across assistants, and whether introduced issues persist or are later addressed. Overall, this work offers new evidence on the maintenance costs behind the rapid adoption of AI coding assistants and highlights the need for stronger quality assurance in AI-assisted software development.